MILLIMETERWAVE SPECTROSCOPY OF THE INTERNAL ROTATION BANDS OF Ne-DCN

NAOKO OYAMADA, \underline{KENSUKE HARADA}, and KEIICHI TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

In 2005,^{a} we reported the MMW specrtum of internal rotation bands (j=1-0 and 2-1) of Ne-HCN to analyzed the intermolecular potential energy surface (PES) between Ne and HCN, where j denotes the quantum number for the HCN internal rotation. In the present study, we have extended our observation to the Ne-DCN deuterated complex in the MMW region (78-175 GHz), and assigned the several DCN internal rotation bands such as the j=1-0 fundamental band (Σ_{1}-Σ_{0} and Π_{1}-Σ_{0}) and the j=2-1 hot band (Σ_{2}-Σ_{1}, Π_{2}-Σ_{1}, Π_{2}-Π_{1}, and Δ_{2}-Π_{1}) for the ^{20}Ne-DCN and ^{22}Ne-DCN complexes. In total, 69 and 12 lines have been assigned to the ^{20}Ne-DCN and ^{22}Ne-DCN. The intermolecular stretch band between Ne and DCN, however, was not observed in this frequency region. Analysis shows that the Σ_{1} and Π_{1} sublevels for j=1 state are located at 134 and 105 GHz, respectively, above the j=0 ground state (Σ_{0}), while the Σ_{2}, Π_{2}, and Δ_{2} sublevels of j=2 state are located at 286, 276, and 257 GHz with different order from that for the normal species. The observed MMW frequencies for Ne-DCN were analyzed with two dimensional (θ − R) PES freezing the freedom in DCN moiety. The PES given by CCSD(T) level ab initio calculation^{b} was modified by adding sixteen extra parameters and fitted to the observed frequencies of internal rotation bands of both ^{20}Ne and ^{22}Ne species. The (θ − R) PES thus fitted has a global minimum in the linear configuration (Ne⋯D–C–N) with a well depth of 64.1 cm^{−1}, and a saddle point located in the anti-linear configuration (D–C–N⋯Ne) by 18.4 cm^{−1} higher than the global minimum. The j=0 ground vibrational state is located by 4.8 cm^{−1} higher than the saddle point. The PES is anisotropic because the center-of-mass distance between Ne and DCN changes much along the minimum energy path, 4.230, 3.477, and 4.020 Å in the linear, T-shaped, and anti-linear forms, together with their energies. The PES estimated for Ne-DCN is very similar to that of Ne-HCN, but the global minimum is by 1.1 cm^{−1} deeper than that of Ne-HCN, due to the frozen model of the HCN/DCN moiety and also our observation is quite limited to the bottom of PES, e.g. highest observed state (Σ_{2}) is still 30 cm^{−1} below the dissociation limit.

aK. Harada, K. Tanaka, and S. Nanbu, The 60th International Symposium on Molecular Spectroscopy, RH01, (2005).

bR. R. Toczylowski, F. Doloresco, and S. M. Cybulski, J. Chem. Phys. 114, 851 (2001).